# Distilling Diffusion Models to Efficient 3D LiDAR Scene Completion

### v1

@@ Abstract
Diffusion models have been applied to 3D LiDAR scene completion due to their strong training stability and high completion quality. However, the slow sampling speed limits the practical application of diffusion-based scene completion models since autonomous vehicles require an efficient perception of surrounding environments. This paper proposes a novel distillation method tailored for 3D LiDAR scene completion models, dubbed ScoreLiDAR, which achieves efficient yet high-quality scene completion. ScoreLiDAR enables the distilled model to sample in significantly fewer steps after distillation. To improve completion quality, we also introduce a novel Structural Loss, which encourages the distilled model to capture the geometric structure of the 3D LiDAR scene. The loss contains a scene-wise term constraining the holistic structure and a point-wise term constraining the key landmark points and their relative configuration. Extensive experiments demonstrate that ScoreLiDAR significantly accelerates the completion time from 30.55 to 5.37 seconds per frame ($>$5$\times$) on SemanticKITTI and achieves superior performance compared to state-of-the-art 3D LiDAR scene completion models. Our code is publicly available on https://github.com/happyw1nd/ScoreLiDAR.
@@ 1 Introduction
Figure 2: A visualization of LiDAR scene completion performances with different models on SemanticKITTI [1] dataset. Our proposed ScoreLiDAR with 8 sampling steps performs better than LiDiff [24] with 50 steps, as shown by a lower Chamfer Distance yet with less time cost. Generally, ScoreLiDAR achieves better scene completion performance and speed trade-off.
Due to the advantages of strong training stability and high-generation quality, existing works utilize diffusion models to complete the 3D LiDAR scenes and achieve outstanding results [24, 23]. However, the diffusion model often requires multiple network iterations to obtain a dense, complete, and high-quality LiDAR scene, which is time-consuming [50, 12]. Autonomous vehicles require fast and efficient perception and recognition of surrounding environments, so the slow sampling speed limits the practical application of diffusion models. Although existing works have proposed acceleration methods for diffusion models [33, 29, 19, 46, 36], 3D LiDAR scenes contain complex geometric structure information, directly applying existing acceleration methods may lead to degraded local details and reduced realism in the completed scene.
In this work, we propose ScoreLiDAR, a novel distillation method tailored for 3D LiDAR scene completion diffusion models, which enables efficient and high-quality scene completion. Variational Score Distillation (VSD) [37] uses a pre-trained diffusion model to calculate a distribution matching loss for training a student model, which has achieved impressive results. Inspired by this, ScoreLiDAR adapts and expands VSD for effective distillation of the pre-trained 3D LiDAR scene completion diffusion model. Moreover, we introduce a Structural Loss to ensure the stability of training and improve the final performance. The structural loss contains a scene-wise term constraining the holistic structure and a point-wise term constraining the key landmark points and their relative configuration, which helps the student model to capture the geometric structure information of 3D LiDAR scenes and achieve high-quality completion. We compared the proposed ScoreLiDAR with the state-of-the-art (SOTA) LiDAR scene completion models. Extensive experiments demonstrate that ScoreLiDAR can effectively accelerate the sampling speed of LiDAR scene completion diffusion models while achieving optimal scene completion quality, as shown in Fig. 1 and Fig. 2.
@@ 2 Related work
3D LiDAR scene completion refers to recovering a complete scene from sparse, incomplete LiDAR scan in applications such as autonomous driving [41, 34]. Current mainstream LiDAR scene completion methods include depth completion-based and Signed Distance Field (SDF)-based approaches. Depth completion-based methods aim to recover dense depth maps from sparse depth measurements [42, 8, 38]. These methods typically leverage deep learning techniques [4, 7] and can also incorporate guidance from RGB images to achieve higher-quality completion results [49, 28, 47]. SDF-based methods represent scenes as voxel grids, with the core idea of using signed distance fields to complete sparse LiDAR scene [34, 14]. These methods are constrained by voxel resolution, making them prone to losing details within the scene [24, 6]. In addition, some methods introduce semantic information to enhance LiDAR scene completion [27, 39]. These methods can generate dense and complete scenes while providing semantic labels for each point, leading to broader application potential [43, 35].
Diffusino-based 3D LiDAR scene completion
Due to the strong training stability and high generation quality of diffusion models, many methods leverage diffusion models for LiDAR scene completion tasks [24, 3, 23, 25, 13]. The work of Lee et al[13] is the first to apply diffusion models at the scene scale for LiDAR scene completion, enabling the generation of realistic scenes conditioned on partial observations from sparse point clouds. Similarly, R2DM [23] utilizes diffusion models based on distance and reflectance intensity image representations to generate various high-fidelity 3D LiDAR scenes. LiDiff [24] indicates that adding noise to point cloud data at the scene scale leads to a loss of detail. Therefore, LiDiff proposes operating directly on individual points and redefines the noise schedule and denoising processes to generate scenes with richer detail. Based on LiDiff, DiffSSC [3] further performs semantic scene completion tasks by implementing denoising and segmentation separately in both the point and semantic spaces. Moreover, LiDMs [25] constructs the pipeline from the perspectives of modal realism, geometric realism, and object realism, achieving generation under different conditions.
However, due to the inherently slow sampling process of diffusion models, the inference of these diffusion-based 3D LiDAR scene completion models is relatively slow. This limitation makes it challenging to achieve fast and efficient perception as required in autonomous vehicle applications.
@@ 3.1 Brief introduction of diffusion models
The diffusion models have two processes: forward diffusion and reverse denoising process [9, 30]. In the forward diffusion process, given the data $\boldsymbol{x}^{0}\sim q(\boldsymbol{x})$ from the training distribution, the diffusion model adds different scales of noise to $\boldsymbol{x}^{0}$ according to different timesteps $t\in[1,T]$ to obtain noisy data $\{\boldsymbol{x}^{1},\boldsymbol{x}^{2},\ldots,\boldsymbol{x}^{T}\}$. When $T$ is large enough, $\boldsymbol{x}^{T}$ approaches to standard Gaussian distribution, namely, $q(x^{T})\approx\mathcal{N}(0,I)$. This process is parameterized by a series of predefined noise factors $\beta^{t}$. By defining $\alpha^{t}=1-\beta^{t}$, the diffusion process is expressed as [9]:
@@
$\mathcal{L}_{DM}=\mathbb{E}_{t,\epsilon}\left[\|\boldsymbol{\epsilon}^{t}-\boldsymbol{\epsilon}_{\theta}(\boldsymbol{x}^{t},c,t)\|^{2}\right]$ (2)
@@
In the reverse denoising process, the diffusion model starts from the timestep $T$ and progressively removes the predicted noise until a generated sample is obtained. The process of denoising $\boldsymbol{x}^{t}$ to obtain $\boldsymbol{x}^{t-1}$ can be written as:
$\boldsymbol{x}^{t-1}=\frac{1}{\sqrt{\alpha^{t}}}\left(\boldsymbol{x}^{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}^{t}}}\boldsymbol{\epsilon}_{\theta}\left(\boldsymbol{x}^{t},c,t\right)\right)+\sigma^{t}\boldsymbol{z}$ (3)
Here $z\sim\mathcal{N}(0,I)$. In this process, the number of required inference steps varies depending on different sampling methods. For instance, DDPM [9] requires 1000 timesteps, while DDIM [30] and DPM solver [18] can reduce this to 100 timesteps and 20 timesteps, respectively.
@@
The 3D LiDAR scene completion diffusion models take the incomplete scan $\mathcal{P}=\{\boldsymbol{p}_{1},\boldsymbol{p}_{2},...,\boldsymbol{p}_{N}\}$ and try to generate the complete scene $\mathcal{G}^{0}=\{\boldsymbol{p}^{0}_{1},\boldsymbol{p}^{0}_{2},...,\boldsymbol{p}^{0}_{M}\}$. Given the input LiDAR scan $\mathcal{P}$ and ground truth $\mathcal{G}$, a diffusion model can be trained to perform 3D LiDAR scene completion. The noisy scene $\mathcal{G}^{t}$ at timestep $t$ is calculated from the ground truth $\mathcal{G}$ at point level [24, 3],
@@
Then, as described in Sec. 3.1, the completed scene ${\mathcal{G}}^{0}$ can be generated by progressive denoising from $\mathcal{G}^{T}$. Because the scale of the LiDAR scene is large and the data range is different across different point cloud axes, directly normalizing the entire dataset compresses the data into a smaller range, which potentially leads to the loss of critical details [24, 3]. To solve this issue, LiDiff [24] modifies the diffusion process by adding a local noise offset to each point $\boldsymbol{p}_{m}$, gradually perturbing the point cloud at each timestep. For Eq. 1, $\boldsymbol{x}_{0}$ is set to 0, and $\boldsymbol{x}_{t}$ is added to each point $\boldsymbol{p}_{m}$,
$\boldsymbol{p}_{m}^{t}=\boldsymbol{p}_{m}+\left(\sqrt{\bar{\alpha}_{t}}\mathbf{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon^{t}}\right)=\boldsymbol{p}_{m}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon^{t}}$ (6)
Due to this special case, $\mathcal{G}^{T}$ cannot directly start from standard Gaussian noise in the sampling process. Instead, the LiDAR scan $\mathcal{P}$ is used to obtain $\mathcal{G}^{T}$ [24]. Firstly, given the initial incomplete scan $\mathcal{P}$, the number of the point clouds is increased by duplicating the original points $K$ times and getting the pseudo dense scan $\mathcal{P}^{*}=\{\boldsymbol{p}^{*}_{1},\boldsymbol{p}^{*}_{2},\ldots,\boldsymbol{p}^{*}_{M}\}$, where we assume $M=KN$. Then, we calculate the noisy point cloud $\mathcal{P}^{T}$ by Eq. 6. As $\mathcal{P}^{T}$ is noisy enough, it can be regarded as $\mathcal{G}^{T}$ during the training. After that, a step-by-step denoising process is applied to obtain the completed scene $\mathcal{G}^{0}$.
@@ 4 Method
Our goal is to distill a pre-trained 3D LiDAR scene completion diffusion model into a student model with significantly fewer sampling steps, enabling efficient and high-quality scene completion. Firstly, we introduce the distillation method tailored for 3D LiDAR scene completion diffusion models in Sec. 4.1. Then, we introduce the structural loss to improve the distillation process with both scene-wise loss and point-wise loss in Sec. 4.2. Finally, we describe the optimization procedure of ScoreLiDAR in Sec. 4.3.
@@ 4.1 Distillation for 3D LiDAR scene completion
$\min_{\eta}\mathcal{L}_{KL}=\mathbb{E}_{t,\epsilon}D_{KL}\left(p^{t}_{G}\left(\mathcal{G}^{t}\right)\|q^{t}\left(\mathcal{G}^{t}\right)\right)$ (8)
@@
We use the pre-trained diffusion model $\boldsymbol{\epsilon}_{\theta}$ to approximate $\nabla_{\mathcal{G}^{t}}\log q^{t}\left(\mathcal{G}^{t}\right)$ with $\nabla_{\mathcal{G}^{t}}\log q^{t}\left(\mathcal{G}^{t}\right)\approx-\frac{\boldsymbol{\hat{\epsilon}}}{\sqrt{1-\bar{\alpha}^{t}}}$ [32]. Similarly, the score $\nabla_{\mathcal{G}^{t}}\log p_{G}^{t}\left(\mathcal{G}^{t}\right)$ can be approximated by another auxiliary diffusion model $\boldsymbol{\epsilon}_{\phi}$. Then, with the simplification as in [9], the $\mathcal{L}_{KL}$ is estimated by
@@
[…]
@@ 4.2 Structural loss
Although the distillation process in Sec. 4.1 is highly effective in training models [19, 50], we found that directly applying it to LiDAR scene completion diffusion models leads to loss of local details and reduced realism in the completed scene. This is because the point cloud in LiDAR scenes often includes complex geometric information that is not explicitly captured by diffusion models. Thus, we introduce a structural loss to further refine the distillation process and improve the completion quality. This structural loss includes scene-wise loss and point-wise loss and can help the student model effectively capture geometric structure information of the 3D point clouds.
@@
To solve this issue, we introduce the scene-wise loss, which minimizes the distance between the ground truth scene $\mathcal{G}$ and the completed scene $\mathcal{G}^{0}$. Concretely, we calculate the scene-wise loss by
$\mathcal{L}_{scene}=\mathbb{E}_{t,\epsilon}\left[\|\mathcal{G}-\mathcal{G}^{0}\|^{2}_{2}\right]$ (12)
This loss calculates the mean squared error between each point in the generated scene and its closest corresponding point in the ground truth. It helps the student model capture the holistic structure, which prevents the optimization direction from deviating in the early stages and enhances training stability. The scene-wise loss also enables the generated scenes to be closer to the ground truth globally, thereby enhancing the completion quality and fidelity.
@@
Given the ground truth $\mathcal{G}=\{\boldsymbol{p}_{1},\boldsymbol{p}_{2},...,\boldsymbol{p}_{M}\}$, we select $n$ key points ($n\ll M$) from the point cloud to construct the $n\times n$ distance matrix $\mathcal{D}$. The $d_{ij}\in\mathcal{D}$ represents the Euclidean distance between the point $i$ and $j$. Then, for completed scene $\mathcal{G}^{0}$, we select $n$ points that are closest to the key points in $\mathcal{G}$ as the corresponding key points to obtain the distance matrix $\mathcal{D}_{G}$. Thus, the point-wise loss is calculated by

### v2

@@ Abstract
Diffusion models have been applied to 3D LiDAR scene completion due to their strong training stability and high completion quality. However, the slow sampling speed limits the practical application of diffusion-based scene completion models since autonomous vehicles require an efficient perception of surrounding environments. This paper proposes a novel distillation method tailored for 3D LiDAR scene completion models, dubbed ScoreLiDAR, which achieves efficient yet high-quality scene completion. ScoreLiDAR enables the distilled model to sample in significantly fewer steps after distillation. To improve completion quality, we also introduce a novel Structural Loss, which encourages the distilled model to capture the geometric structure of the 3D LiDAR scene. The loss contains a scene-wise term constraining the holistic structure and a point-wise term constraining the key landmark points and their relative configuration. Extensive experiments demonstrate that ScoreLiDAR significantly accelerates the completion time from 30.55 to 5.37 seconds per frame ($>$5$\times$) on SemanticKITTI and achieves superior performance compared to state-of-the-art 3D LiDAR scene completion models. Our model and code are publicly available on https://github.com/happyw1nd/ScoreLiDAR.
@@ 1 Introduction
Figure 2: A visualization of LiDAR scene completion performances with different models on SemanticKITTI [1] dataset. Generally, our proposed ScoreLiDAR achieves better scene completion performance and speed trade-off.
Due to the advantages of strong training stability and high-generation quality, existing works utilize diffusion models to complete the 3D LiDAR scenes and achieve outstanding results [28, 27]. However, the diffusion model often requires multiple network iterations to obtain a dense, complete, and high-quality LiDAR scene, which is time-consuming [54, 13]. Autonomous vehicles require fast and efficient perception and recognition of surrounding environments, so the slow sampling speed limits the practical application of diffusion models. Although existing works have proposed acceleration methods for diffusion models [37, 33, 22, 49, 40], due to the differences between 3D LiDAR scenes and image data—where LiDAR scenes often contain complex geometric structure information—these techniques have not been explored in the acceleration of diffusion-based LiDAR scene completion model.
In this work, we propose ScoreLiDAR, a novel distillation method tailored for 3D LiDAR scene completion diffusion models, which enables efficient and high-quality scene completion (Fig. 1 and Fig. 2). ScoreLiDAR aims to tackle the unique 3D distribution alignment challenge in LiDAR scene completion. By exploiting a bidirectional gradient guidance mechanism, it allows the student model’s completion results under sparse point cloud conditions to progressively approach the multi-step iterative reconstruction quality of the teacher model. Given a completed scene generated by the student model, the teacher model predicts the scene’s score as a gradient to enhance realism, while a designed auxiliary model predicts the scene’s another score as a gradient to suppress unrealistic completion. The student model is then updated according to the difference between these two gradients. This update drives the completion process toward more structurally coherent LiDAR scenes. Finally, we introduce a Structural Loss consisting of a scene-wise term and a point-wise term constraining the key landmark points and their relative configuration. Prior studies [42, 41, 23, 19] demonstrated that the bidirectional gradient guidance mechanism can effectively accelerate 3D rendering speed. In this work, we argue that a similar approach can be directly applied to the distillation of diffusion-based LiDAR scene completion models.
@@ 2 Related work
3D LiDAR scene completion refers to recovering a complete scene from a sparse, incomplete LiDAR scan in applications such as autonomous driving [46, 38]. Current mainstream LiDAR scene completion methods include depth completion-based and Signed Distance Field (SDF)-based approaches. Depth completion-based methods aim to recover dense depth maps from sparse depth measurements [47, 9, 43]. These methods typically leverage deep learning techniques [4, 7] and can also incorporate guidance from RGB images to achieve higher-quality completion results [53, 32, 50]. SDF-based methods represent scenes as voxel grids, with the core idea of using signed distance fields to complete sparse LiDAR scenes [38, 16]. These methods are constrained by voxel resolution, making them prone to losing details within the scene [28, 6]. In addition, some methods introduce semantic information to enhance LiDAR scene completion [31, 44]. These methods can generate dense and complete scenes while providing semantic labels for each point, leading to broader application potential [48, 39].
Diffusion-based 3D LiDAR scene completion
Due to the strong training stability and high generation quality of diffusion models, many methods leverage diffusion models for LiDAR scene completion tasks [28, 3, 27, 29, 14]. The work of Lee et al[14] is the first to apply diffusion models at the scene scale for LiDAR scene completion, enabling the generation of realistic scenes conditioned on partial observations from sparse point clouds. Similarly, R2DM [27] utilizes diffusion models based on distance and reflectance intensity image representations to generate various high-fidelity 3D LiDAR scenes. LiDiff [28] indicates that adding noise to point cloud data at the scene scale leads to a loss of detail. Therefore, LiDiff proposes operating directly on individual points and redefines the noise schedule and denoising processes to generate scenes with richer detail. Based on LiDiff, DiffSSC [3] further performs semantic scene completion tasks by implementing denoising and segmentation separately in both the point and semantic spaces. Moreover, LiDMs [29] constructs the pipeline from the perspectives of pattern realism, geometric realism, and object realism, achieving generation under different conditions.
Due to the slow sampling inherent in diffusion models, diffusion-based 3D LiDAR scene completion suffers from slow inference, hindering its application in autonomous vehicles. Thus, this paper proposes a distillation method for diffusion-based 3D LiDAR scene completion models to achieve faster and higher-quality completion.
@@ 3.1 Brief introduction of diffusion models
The diffusion models have two processes: forward diffusion and reverse denoising process [10, 34]. In the forward diffusion process, given the data $\boldsymbol{x}^{0}\sim q(\boldsymbol{x})$ from the training distribution, the diffusion model adds different scales of noise to $\boldsymbol{x}^{0}$ according to different timesteps $t\in[1,T]$ to obtain noisy data $\{\boldsymbol{x}^{1},\boldsymbol{x}^{2},\ldots,\boldsymbol{x}^{T}\}$. When $T$ is large enough, $\boldsymbol{x}^{T}$ approaches to standard Gaussian distribution, namely, $q(\boldsymbol{x}^{T})\approx\mathcal{N}(0,I)$. This process is parameterized by a series of predefined noise factors $\beta^{t}$. By defining $\alpha^{t}=1-\beta^{t}$, the diffusion process is expressed as [10]:
@@
$\mathcal{L}_{DM}=\mathbb{E}_{t,\epsilon}\left[\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(\boldsymbol{x}^{t},c,t)\|^{2}\right]$ (2)
@@
In the reverse denoising process, the diffusion model starts from the timestep $T$ and progressively removes the predicted noise to obtain a generated sample. The process of denoising $\boldsymbol{x}^{t}$ to obtain $\boldsymbol{x}^{t-1}$ can be written as in [10]
$\boldsymbol{x}^{t-1}=\frac{1}{\sqrt{\alpha^{t}}}\left(\boldsymbol{x}^{t}-\frac{1-\alpha^{t}}{\sqrt{1-\bar{\alpha}^{t}}}\boldsymbol{\epsilon}_{\theta}\left(\boldsymbol{x}^{t},c,t\right)\right)+\sigma^{t}\textbf{z}$ (3)
Here $\textbf{z}\sim\mathcal{N}(0,I)$. In this process, the number of required inference steps varies depending on different sampling methods. For instance, DDPM [10] requires 1000 timesteps, while DDIM [34] and DPM solver [21] can reduce this to no more than 100 timesteps.
@@
The 3D LiDAR scene completion diffusion models take the incomplete scan $\mathcal{P}=\{\boldsymbol{p}_{1},\boldsymbol{p}_{2},...,\boldsymbol{p}_{N}\}$ ( $p_{i}\in\mathbb{R}^{3}$) and try to generate the complete scene $\mathcal{G}^{0}=\{\boldsymbol{p}^{0}_{1},\boldsymbol{p}^{0}_{2},...,\boldsymbol{p}^{0}_{M}\}$. Given the input LiDAR scan $\mathcal{P}$ and ground truth $\mathcal{G}$, a diffusion model can be trained to perform 3D LiDAR scene completion. The noisy scene $\mathcal{G}^{t}$ at timestep $t$ is calculated from the ground truth $\mathcal{G}$ at point level [28, 3],
@@
Then, as described in Sec. 3.1, the completed scene ${\mathcal{G}}^{0}$ can be generated by progressive denoising from $\mathcal{G}^{T}$. Because the scale of the LiDAR scene is large and the data range is different across different point cloud axes, directly normalizing the entire scene compresses the data into a smaller range, which potentially leads to the loss of critical details [28, 3]. To solve this issue, LiDiff [28] modifies the diffusion process by adding a local noise offset to each point $\boldsymbol{p}_{m}$, gradually perturbing the point cloud at each timestep. For Eq. 1, $\boldsymbol{x}^{0}$ is set to 0, and $\boldsymbol{x}^{t}$ is added to each point $\boldsymbol{p}_{m}$,
$\boldsymbol{p}_{m}^{t}=\boldsymbol{p}_{m}+\left(\sqrt{\bar{\alpha}^{t}}\mathbf{0}+\sqrt{1-\bar{\alpha}^{t}}\boldsymbol{\epsilon^{t}}\right)=\boldsymbol{p}_{m}+\sqrt{1-\bar{\alpha}^{t}}\boldsymbol{\epsilon^{t}}$ (6)
@@
Due to this special case, the initial noisy scene $\mathcal{G}^{T}$ cannot directly start from standard Gaussian noise in the sampling process. Instead, the LiDAR scan $\mathcal{P}$ is used to obtain $\mathcal{G}^{T}$ [28]. Firstly, given the initial incomplete scan $\mathcal{P}$, the number of the point clouds is increased by duplicating the original points $K$ times and getting the pseudo dense scan $\mathcal{P}^{*}=\{\boldsymbol{p}^{*}_{1},\boldsymbol{p}^{*}_{2},\ldots,\boldsymbol{p}^{*}_{M}\}$, where we assume $M=KN$. Then, we calculate the noisy point cloud $\mathcal{P}^{T}$ by Eq. 6. As $\mathcal{P}^{T}$ is noisy enough, it can be regarded as $\mathcal{G}^{T}$ during the training. After that, a step-by-step denoising process is applied to obtain the completed scene $\mathcal{G}^{0}$.
@@ 4 Method
Our goal is to distill a pre-trained 3D LiDAR scene completion diffusion model into a student model with significantly fewer sampling steps, enabling efficient and high-quality scene completion. Firstly, we introduce the distillation method tailored for 3D LiDAR scene completion diffusion models in Sec. 4.1. Then, we introduce the structural loss to improve the distillation process with both scene-wise loss and point-wise loss in Sec. 4.2. Finally, we describe the optimization procedure of ScoreLiDAR in Sec. 4.3. The overall structure of ScoreLiDAR is shown in Fig. 3.
@@ 4.1 Distillation for 3D LiDAR scene completion
$\min_{\eta}\mathcal{L}_{KL}=\mathbb{E}_{t,\epsilon}\left[D_{KL}\left(p^{t}_{G}\left(\mathcal{G}^{t}\right)\|q^{t}\left(\mathcal{G}^{t}\right)\right)\right]$ (8)
@@
We use the pre-trained diffusion model $\boldsymbol{\epsilon}_{\theta}$ to approximate $-\sqrt{1-\bar{\alpha}^{t}}\nabla_{\mathcal{G}^{t}}\log q^{t}\left(\mathcal{G}^{t}\right)$ as discussed in Section 5 of Supplementary Materials (Sec S5). Similarly, $-\sqrt{1-\bar{\alpha}^{t}}\nabla_{\mathcal{G}^{t}}\log p_{G}^{t}\left(\mathcal{G}^{t}\right)$ is approximated by an auxiliary diffusion model $\boldsymbol{\epsilon}_{\phi}$, which is independently trained with the denoising loss Eq. 5 on generated samples $\mathcal{G}^{0}$. Then, with the simplification as in [10, 42], the $\mathcal{L}_{KL}$ is estimated by
@@
Intuitively, the orientation of $-\boldsymbol{\epsilon}_{\theta}(\mathcal{G}^{t},\mathcal{P},t)$ points to the pre-trained distribution, while that of $-\boldsymbol{\epsilon}_{\phi}(\mathcal{G}^{t},\mathcal{P},t)$ points to the student model’s generative distribution. Thus, decenting along the bidirectional gradient $\left[\boldsymbol{\epsilon}_{\theta}\left(\mathcal{G}^{t},\mathcal{P},t\right)-\boldsymbol{\epsilon}_{\phi}\left(\mathcal{G}^{t},\mathcal{P},t\right)\right]$ updates the student model’s generative distribution toward the pre-trained distribution, achieving a more accurate completion. Sec S2.1 discusses the efficiency of the distillation process.
@@ 4.2 Structural loss
Although the distillation process in Sec. 4.1 is effective in training models [22, 54], we found that directly applying it to LiDAR scene completion diffusion models leads to loss of local details and reduced realism. This is because the point cloud in LiDAR scenes includes complex geometric information that is not explicitly captured by diffusion models. Thus, we introduce a structural loss to further refine the distillation process and improve the completion quality. This structural loss includes scene-wise loss and point-wise loss and can help the student model effectively capture geometric structure information of the 3D point clouds.
@@
To solve this issue, we introduce the scene-wise loss, which minimizes the distance between the ground truth scene $\mathcal{G}$ and the completed scene $\mathcal{G}^{0}$,
$\mathcal{L}_{scene}=\frac{1}{|\mathcal{G}^{0}|}\sum_{\boldsymbol{p}^{0}_{i}\in\mathcal{G}^{0}}\min_{\boldsymbol{p}\in\mathcal{G}}\|\boldsymbol{p}^{0}_{i}-\boldsymbol{p}\|^{2}$ (12)
This loss calculates the mean squared error between each point $\boldsymbol{p}^{0}_{i}$ in the generated scene $\mathcal{G}^{0}$ and its closest corresponding point $\boldsymbol{p}$ in the ground truth $\mathcal{G}$. It helps the student model capture the holistic structure, which prevents the optimization direction from deviating in the early stages and enhances training stability. The scene-wise loss enables the generated scenes to be closer to the ground truth globally, thereby enhancing the completion quality and fidelity.
@@
Given the ground truth $\mathcal{G}=\{\boldsymbol{p}_{1},\boldsymbol{p}_{2},\ldots,\boldsymbol{p}_{M}\}$, we select $n$ key points ($n\ll M$) from the point cloud to construct the $n\times n$ distance matrix $\mathcal{D}$. The $d_{ij}\in\mathcal{D}$ represents the pairwise Euclidean Distance between points $i$ and $j$. Then, for completed scene $\mathcal{G}^{0}$, we select $n$ points that are closest to the key points in $\mathcal{G}$ as the corresponding key points to obtain the distance matrix $\mathcal{D}_{G}$. Thus, the point-wise loss is calculated by